ddcdcc:

Why is x = 4 a solution to the proportion StartFraction 14 over x EndFraction = StartFraction 56 over 16 EndFraction?

Mercury:

\[\frac{ 14 }{ x } = \frac{ 56 }{ 16 } \] we are told that x = 4 is a solution to this equation. therefore, when you plug in 4 where x is, both sides should still be equal. \[\frac{ 14 }{ 4 } = \frac{ 56 }{ 16 } \] there's a couple ways you could prove this: - cross multiply and verify that both sides are equal - find a constant that you can multiply the left side by, to get the right side - convert both sides to a decimal by dividing the numerator by the denominator take your pick